OpenStudy (anonymous):

f(x)=1/x+2 reflected across the y-axis

OpenStudy (anonymous):

1/(x+2) x+2 is the denominator

OpenStudy (anonymous):

no it 1/x plus 2

OpenStudy (anonymous):

f(x) = 1/x + 2 The equation of the reflection of f(x) about the y-axis is y = f(−x). f(-x) = 1/(-x) +2 = -1/x + 2

OpenStudy (anonymous):

i meant vertical compressin sorry

OpenStudy (anonymous):

Reflection about x axis

OpenStudy (anonymous):

y = f(x). Its reflection about the x-axis is y = −f(x). f(x) = 1/x +2 -f(x) = -( 1/x +2) = -1/x -2 is the reflection about x axis